# Studies of stationary features in jets: 3C 279 quasar I. On-sky scattering and dynamics

### v1

@@ 5 Characteristics of relativistic transverse waves
In this section, we analyse the QSC trajectory to evaluate the transverse wave frequency and amplitude. The observed trajectory of the QSC is composed of apparent displacements, $r$, between successive positions of the QSC. The apparent displacement is a combination of the intrinsic apparent displacement of the QSC ($s$) and the displacement of the core ($c$). To recover the intrinsic trajectory of the QSC, it is necessary to smooth out the anisotropic displacements of the core, which occur along the jet axis. To achieve this, we apply a moving average procedure to QSC positions, thereby mitigating the effect of core displacement. In this method, a time window centered at each epoch rolls over the epochs and calculates the average position and its associated error within the smoothing time window ($\Delta t_{\rm opt}$). Edge treatment: the method excludes epochs from the moving average calculation when the full time window cannot be applied.
The length of smoothing window.
@@
The time interval for the sliding window should be chosen optimally to average out the effects of core displacement while preserving, as much as possible, the intrinsic trajectories of QSC. The length of time window should be large enough to smooth out wiggling of the core due to physical reasoning such as particle density and magnetic field changes which may result by passage of moving plasma perturbations through the core (Arshakian et al. 2024).
@@
Plavin et al. (2019) suggested that moving components, such as shocks or plasma blobs, can influence the observed core position and brightness. These moving components can lead to temporary increases in core brightness and shift the apparent core position upstream when approaching the core and downstream, in direction of its motion, when moving away from the core. In the same way, moving components can cause apparent displacements of the QSC position when they pass through the QSC region. This will lead to correlated displacements of the core and QSC with a time delay equal to the time that the perturbation spent traveling the distance between the core and QSC. To estimate the time lag we applied the z-transformed discrete cross-correlation function (zDCF; Alexander 1997) to variation of the flux density of the core and QSC with epoch. We found that the maximum correlation $\tau=0.42$ reaches at a time lag of about 0.7 yr, where the variation of the core flux density is leading (Fig. 10). We assume that the length of the sliding time window should be $\gtrsim 0.7$ years to properly smooth the correlated displacements of the core and QSC.
@@
Another issue is how often the disturbance passes through the core, i.e., the ejection rate of moving components. The quasar 3C 279 has been extensively studied to determine the ejection rate of moving components. Jorstad et al. (2001) monitored 3C 279 with VLBA at 22 GHz and 43 GHz over several years and identified multiple superluminal components. They estimated an average ejection rate of approximately one new component per year and reported variability in the ejection rate, with periods of increased activity where multiple components were ejected within a single year, followed by quieter periods with fewer or no ejections. We assume that the length of the smoothing window should be equal or longer than the timescales of the core disturbances due to moving components $\Delta t_{\rm opt}\gtrsim 1$ yr.
@@
Arshakian et al. (2024) noted that another possible bias in the fitted position of the components, potentially leading to apparent regular variations in the position of QSC in the BL Lac jet, may arise from the model fitting process. This is not the case with 3C 279 since the core is very bright and compact with a relatively high median cadence of observations of about 40 days.
@@
Finally, we conduct a decision analysis of the QSC position displacements along and across the jet axis (black lines in Fig. 11) to determine the optimal width of the smoothing window. The variations along the jet axis are more pronounced due to the core’s oscillatory displacements along the jet direction. To effectively smooth these variations, the sliding window length must be sufficiently large to capture long-term variations on the order of one year (bottom panel). We adopt $\Delta t_{\rm opt}=1.5$ yr, which effectively smooths variations along and across the jet axis on timescales of one year or longer, as illustrated by the blue lines in Fig. 11.
@@
Figure 12: The trajectory of the QSC from 1995.57 to 2022.22, smoothed using a time window of 1.5 years (blue line). The dashed light blue line connects the endpoints of the observation gap. Red ellipses indicate the average asymmetric positioning uncertainties. The median scatter position is marked with a red plus sign, while the dashed red line represents the central jet axis, connecting the median QSC position to the radio core.
@@
The trajectory smoothed by moving average procedure on timescales of 1.5 yr is presented in Fig. 12. The number of averaged positions is reduced to 139 due to boundary handling at the epoch endpoints and a gap in observations. Each averaged position is associated with averaged asymmetric position error, which is shown as an ellipse, with its major axis oriented toward the core. The asymmetric $1\sigma$ position errors are represented by half the lengths of the major and minor axes. The mean position errors are relatively small and generally do not overlap, except in regions where the QSC positions exhibit clustering. Following Arshakian et al. (2024), we consider displacement measurement to be unreliable if the error ellipses of two consecutive positions intersect along the displacement vector. We adopted their algorithm based on this criterion and applied it to the averaged positions. For clarity, we present the refined QSC trajectory in two time intervals (Fig. 13). This procedure further reduces the number of averaged positions by 59%, resulting in a total of 68.
The overall refined trajectory consists of quasi-linear or arc-shaped patterns followed by loop-like patterns or sharp turns, closely resembling the motion of QSC C7 in the BL Lac jet (Arshakian et al. 2024). To determine whether the refined trajectory deviates from randomness, we analyze the angles between successive displacement vectors. In the case of purely random motion, the ratio of the number of angles greater than $90^{\circ}$ to those less than $90^{\circ}$ should be approximately unity, that is, $n(>90^{\circ})/n(<90^{\circ})\approx 1$. Our estimate yields a ratio of $42/13\approx 3.2$, indicating that forward-directed displacement vectors outnumber backward-directed ones by a factor of 3.2. This evidence indicates that the trajectory is not random but follows a directed path with subsequent reversals.
5.1 Characteristics of reversals.
We apply the formalism described in Arshakian et al. (2024) to characterize the reverse motion of the QSC. We classify the trajectory as reversible (R-type) if the turning angle is less than $90\degr$ and combo reversal (RC-type) if a reversal consists of multiple sub-reversals (Table 2), for example, two pivot points are encompassed within the loop-like trajectory spanning 1998.41--2000.44 (Fig. 13, yellow line, left panel).
@@
The direction of reversal at the pivot point is clockwise (CW) in 8 cases and counterclockwise (CCW) in 3 cases (Table 2). Notably, eight consecutive reversals (numbered 3 to 10) follow a CW direction over a 15.52-year period (excluding a 2.36-year observation gap). The probability of this occurring by chance is very small, only 0.4%.
We analysed the on-sky distribution of pivot points, specifically examining their radial distances from the median center and their azimuthal angles relative to the jet axis. The azimuthal angle distribution of the pivot points appears uniform, while most radial distances are concentrated within 0.1 mas from the median center. No significant differences are found in these distributions between CW and CCW motion.
Figure 13: Refined trajectory of the QSC represented for two time periods: 1995.57--2005.71 and 2005.71--2022.22. The arrows denote the direction of movement, and the red circles mark reversal point. The colored segments of the trajectory represent the path lengths of the identified reversals. In the right panel, the black line connects the endpoints of the observation gap. The median scatter position over the whole time range is marked with a red plus sign, while the dashed red line represents the central jet axis, connecting the median QSC position to the radio core.
@@
Arshakian et al. (2024) demonstrate that, in the seagull-on-wave model, the reversing trajectories of the QSC are the on-sky projections of its motion, induced by the passage of a relativistic transverse wave. In this scenario, the amplitude and frequency (or wavelength) of the wave approximately correspond to the frequency and half-length of the reversal, with $f_{\rm w}=f_{\rm R}\sim 0.5$ yr-1, $A_{\rm w}\approx\Delta l_{\rm R}/2\sim 0.06$ mas and $\sigma_{\rm A_{\rm w}}=0.03$ mas. The speeds with which the QSC passes the reversal patterns are superluminal (Table 2). The superluminal speeds can also be explained within the framework of the seagull-on-wave model, where the QSC moves at a subluminal speed in the host galaxy’s rest frame, but appears superluminal in the observer’s frame due to the relativistic motion of the transverse wave. In reality, these speeds represent lower limits, as the QSC trajectory is shortened due to smoothing effects.
@@
Table 2: Characteristics of reversals and reversal combinations. The latter are marked with a gray bar.
@@
Meanc: -- -- -- -- $1.87\pm 0.26$ $0.12\pm 0.01$ $4.36\pm 0.33$ $0.07\pm 0.01$-
 [^3] The columns for reversals are as follows: Reversal type, the epochs of the beginning and end of the reversal, $t_{\rm s}$ and $t_{\rm e}$, reversal time interval $\Delta t_{\rm R}=t_{\rm e}-t_{\rm s}$, length of the reversal $\Delta l_{\rm R}$, speed of the QSC component $\beta^{*}_{\rm s}$ and spatial scale which represents the largest size of the reversal.
@@ 6 Discussions
BL Lac objects and flat-spectrum radio quasars (FSRQ) differ significantly in several key characteristics, including optical spectral properties, jet power, large-scale radio morphology (FR I vs. FR II), spectral energy distributions (synchrotron-dominated vs. inverse Compton-dominated), and the relative strength of their $\gamma$-ray emission. Here, we compare the properties of relativistic transverse waves observed in two blazars: BL Lac (Cohen et al. 2015; Arshakian et al. 2024) and the FSRQ 3C 279, as analysed in this work. The wave frequency and mean wave amplitude are estimated to be $0.5$ yr-1 and 0.06 mas (0.38 pc) for the 3C 279 and $0.8$ yr-1 and 0.028 mas (0.037 pc) for the BL Lac. While the wave frequencies are comparable, the mean wave amplitude of the 3C 279 is larger than that of the BL Lac by a factor of ten. Assuming that the transverse wave is an Alfvén wave, and the transverse magnetic energy $B_{\perp}$ dominates over the transverse kinetic energy of the jet plasma, the total energy of the wave in the frame of the jet $E_{\rm tot}\propto B_{\perp}^{2}V\approx(A_{\rm w}f_{\rm w}B_{0})^{2}V$, where the magnetic energy of the perturbation $B_{\perp}\approx A_{\rm w}f_{\rm w}B_{0}$, $B_{0}$ is the mean large-scale magnetic field that exists before perturbations by Alfvén waves, and $V$ is the total volume of the jet segment containing the wave. Assuming that the values of $B_{0}$ and $V$ are approximately equal for BL Lac and 3C 279, we obtain the ratio of their total energies $(0.5\times 0.38)^{2}/(0.8\times 0.037)^{2}\approx 40$, that is, the total transverse wave energy of 3C 279 is approximately $10^{1.6}$ times higher than that of the BL Lac wave. This difference is primarily due to the significantly larger mean wave amplitude of the quasar 3C 279.
@@
The mean speeds of the QSCs in 3C 279 and BL Lac are superluminal, $15c$ and $2.5c$ (Arshakian et al. 2024), with the former being about six times higher. This difference can be attributed to a smaller viewing angle of the jet, a higher inclination angle and speed of the transverse wave in 3C 279, or a combination of these factors (see Eq. 1).
@@
Another significant difference lies in the jet structure on pc-scales: the deprojected distance of the QSC from the core in 3C 279 ($\approx 125$ pc) is approximately 40 times greater than that of the QSC in BL Lac ($\approx 3$ pc), assuming a jet viewing angle of $1\degr$ adopted in this work and $6\degr$ from Arshakian et al. (2024), respectively. The QSC in the BL Lac jet is believed to be a recollimation shock at 15 GHz (Cohen et al. 2014), supported by high-resolution observations at 22 GHz with RadioAstron and 43 GHz with VLBA (Gómez et al. 2016). Relativistic magnetohydrodynamics simulations in Gómez et al. (2016) show that recollimation shocks appear at relatively short distances from the core due to the jet’s high speed and a decreasing pressure gradient in the ambient medium, possibly caused by interactions with the broad-line region ($0.1-1$ pc) or torus ($\lesssim 10$ pc). On the other hand, the location of the QSC in 3C 279 at a distance of about 125 pc from the core is beyond the broad-line region and torus, indicating that the recollimation shock is generated by interaction with a hot gas region.
@@
Burd et al. (2022) show that the transition of the jet geometry from parabolic to conical in 3C 279 occurs at a separation of 0.6 mas (deprojected distance of $\approx 210$ pc), likely marking the transition from a magnetically dominated to a particle-dominated jet (Kovalev et al. 2020). In BL Lac, the break point of the jet geometry transition occurs at a separation of 2.5 mas (deprojected distance of $\approx 30$ pc; Kovalev et al. 2020), well beyond the torus region. This supports the dominance of helical magnetic energy over kinetic energy of the jet at large distances from the central black hole, and consequently the presence of relativistic Alfvén waves extending up to the jet geometry transition location.
@@
Lastly, the ratio of rmsc/rmss is $\approx 1.9$ for 3C 279 and $\approx 1.3$ for BL Lac (Arshakian et al. 2020), indicating that the variation in the displacements of the core is more pronounced in 3C 279. This could be attributed to stronger opacity changes (or higher particle density variations) during radio flares or a stronger magnetic field in the core region, as well as geometric effects such as wobbling of the core (see a detailed discussion in Arshakian et al. 2020).

### v2

@@ 5 Characteristics of relativistic transverse waves
In this section, we analyse the QSC trajectory to evaluate the transverse wave frequency and amplitude. The observed trajectory of the QSC is composed of apparent displacements, $r$, between successive positions of the QSC. The apparent displacement is a combination of the intrinsic apparent displacement of the QSC ($s$) and the displacement of the core ($c$). To recover the intrinsic trajectory of the QSC, it is necessary to smooth out the anisotropic displacements of the core, which occur along the jet axis. To achieve this, we apply a moving average procedure to QSC positions, thereby mitigating the effect of core displacement. In this method, a time window centred on each epoch rolls over the epochs and calculates the average position and its associated error within the smoothing time window ($\Delta t_{\rm opt}$). Edge treatment: the method excludes epochs from the moving average calculation when the full time window cannot be applied.
The length of the smoothing window.
@@
The time interval for the sliding window should be chosen optimally to average out the effects of core displacement while preserving, as much as possible, the intrinsic trajectories of QSC. The length of time window should be large enough to smooth out wiggling of the core due to physical reasoning such as particle density and magnetic field changes that may result from the passage of moving plasma perturbations through the core (arshakian24).
@@
plavin19 suggested that moving components, such as shocks or plasma blobs, can influence the observed core position and brightness. These moving components can lead to temporary increases in core brightness and shift the apparent core position upstream when approaching the core and downstream, in direction of its motion, when moving away from the core. In the same way, moving components can cause apparent displacements of the QSC position when they pass through the QSC region. This will lead to correlated displacements of the core and QSC with a time delay equal to the time that the perturbation spent travelling the distance between the core and QSC. To estimate the time lag we applied the z-transformed discrete cross-correlation function (zDCF; alexander97) to variation in the flux density of the core and QSC with epoch. We found that the maximum correlation $\tau=0.42$ reaches at a time lag of about 0.7 yr, where the variation in the core flux density is leading (Fig. 9). We assume that the length of the sliding time window should be $\gtrsim 0.7$ years to properly smooth the correlated displacements of the core and QSC.
@@
Another issue is how often the disturbance passes through the core, i.e. the ejection rate of moving components. The quasar 3C 279 has been extensively studied to determine the ejection rate of moving components. jorstad01 monitored 3C 279 with VLBA at 22 GHz and 43 GHz over several years and identified multiple superluminal components. They estimated an average ejection rate of approximately one new component per year and reported variability in the ejection rate, with periods of increased activity where multiple components were ejected within a single year, followed by quieter periods with fewer or no ejections. We assume that the length of the smoothing window should be equal or longer than the timescales of the core disturbances due to moving components $\Delta t_{\rm opt}\gtrsim 1$ yr.
@@
arshakian24 noted that another possible bias in the fitted position of the components, potentially leading to apparent regular variations in the position of QSC in the BL Lac jet, may arise from the model fitting process. This is not the case with 3C 279 since the core is very bright and compact with a relatively high median cadence of observations of about 40 days.
@@
Finally, we conducted a decision analysis of the QSC position displacements along and across the jet axis (black lines in Fig. 10) to determine the optimal width of the smoothing window. The variations along the jet axis are more pronounced due to the core’s oscillatory displacements in the direction of the jet. To effectively smooth these variations, the sliding window length must be sufficiently large to capture long-term variations on the order of one year (bottom panel). We adopt $\Delta t_{\rm opt}=1.5$ yr, which effectively smooths variations along and across the jet axis on timescales of one year or longer, as illustrated by the blue lines in Fig. 10.
@@
Figure 11: Trajectory of the QSC from 1995.57 to 2022.22, smoothed using a time window of 1.5 years (blue line). The dashed light blue line connects the endpoints of the observation gap. Red ellipses indicate the average asymmetric positioning uncertainties. The median scatter position is marked with a red plus sign, while the dashed red line represents the central jet axis, connecting the median QSC position to the radio core.
@@
The trajectory smoothed by moving average procedure on timescales of 1.5 yr is presented in Fig. 11. The number of averaged positions is reduced to 139 due to boundary handling at the epoch endpoints and a gap in observations. Each averaged position is associated with averaged asymmetric position error, which is shown as an ellipse, with its major axis oriented towards the core. The asymmetric $1\sigma$ position errors are represented by half the lengths of the major and minor axes. The mean position errors are relatively small and generally do not overlap, except in regions where the QSC positions exhibit clustering. Following Arshakian et al. (2024), we consider displacement measurement to be unreliable if the error ellipses of two consecutive positions intersect along the displacement vector. We adopted their algorithm based on this criterion and applied it to the averaged positions. For clarity, we present the refined QSC trajectory in two time intervals (Fig. 12). This procedure further reduces the number of averaged positions by 59%, resulting in a total of 68.
The overall refined trajectory consists of quasi-linear or arc-shaped patterns followed by loop-like patterns or sharp turns, closely resembling the motion of QSC C7 in the BL Lac jet (arshakian24). To determine whether the refined trajectory deviates from randomness, we analyse the angles between successive displacement vectors. In the case of purely random motion, the ratio of the number of angles greater than $90^{\circ}$ to those less than $90^{\circ}$ should be approximately unity; that is, $n(>90^{\circ})/n(<90^{\circ})\approx 1$. Our estimate yields a ratio of $42/13\approx 3.2$, indicating that forward-directed displacement vectors outnumber backward-directed ones by a factor of 3.2. This evidence indicates that the trajectory is not random but follows a directed path with subsequent reversals.
Characteristics of reversals.
We apply the formalism described in arshakian24 to characterize the reverse motion of the QSC. We classify the trajectory as reversible (R-type) if the turning angle is less than $90\degr$ and combo reversal (RC-type) if a reversal consists of multiple sub-reversals (Table 2), for example, two pivot points are encompassed within the loop-like trajectory spanning 1998.41-2000.44 (Fig. 12, yellow line, left panel).
@@
The direction of reversal at the pivot point is clockwise (CW) in 8 cases and counter-clockwise (CCW) in 3 cases (Table 2). Notably, eight consecutive reversals (numbered 3 to 10) follow a CW direction over a 15.52-year period (excluding a 2.36-year observation gap). The probability of this occurring by chance is very small, only 0.4%.
We analysed the on-sky distribution of pivot points, specifically examining their radial distances from the median centre and their azimuthal angles relative to the jet axis. The azimuthal angle distribution of the pivot points appears uniform, while most radial distances are concentrated within 0.1 mas from the median centre. No significant differences are found in these distributions between CW and CCW motion.
Figure 12: Refined trajectory of the QSC represented for two time periods: 1995.57-2005.71 and 2005.71-2022.22. The arrows denote the direction of movement, and the red circles mark reversal point. The coloured segments of the trajectory represent the path lengths of the identified reversals. In the right panel, the black line connects the endpoints of the observation gap. The median scatter position over the whole time range is marked with a red plus sign, while the dashed red line represents the central jet axis, connecting the median QSC position to the radio core.
@@
arshakian24 demonstrate that, in the seagull-on-wave model, the reversing trajectories of the QSC are the on-sky projections of its motion, induced by the passage of a relativistic transverse wave. In this scenario, the amplitude and frequency (or wavelength) of the wave approximately correspond to the frequency and half-length of the reversal, with $f_{\rm w}=f_{\rm R}\sim 0.5$ yr-1, $A_{\rm w}\approx\Delta l_{\rm R}/2\sim 0.06$ mas and $\sigma_{\rm A_{\rm w}}=0.03$ mas. The speeds with which the QSC passes the reversal patterns are superluminal (Table 2). The superluminal speeds can also be explained within the framework of the seagull-on-wave model, where the QSC moves at a subluminal speed in the host galaxy’s rest frame, but appears superluminal in the observer’s frame due to the relativistic motion of the transverse wave. In reality, these speeds represent lower limits, as the QSC trajectory is shortened due to smoothing effects.
@@
Table 2: Characteristics of reversals and reversal combinations. [^3] The columns for reversals are as follows: reversal type; the epochs of the beginning and end of the reversal, $t_{\rm s}$ and $t_{\rm e}$; reversal time interval $\Delta t_{\rm R}=t_{\rm e}-t_{\rm s}$; length of the reversal $\Delta l_{\rm R}$; speed of the QSC component $\beta_{\rm s}$; spatial scale, representing the largest extent of the reversal; and the direction of the reverse motion.
@@
Meanc: - - - - $1.87\pm 0.26$ $0.12\pm 0.01$ $4.36\pm 0.33$ $0.07\pm 0.01$ -
@@ 6 Discussions
BL Lac objects and FSRQs differ significantly in several key characteristics, including optical spectral properties, jet power, large-scale radio morphology (FR I vs. FR II), spectral energy distributions (synchrotron-dominated vs. inverse Compton-dominated), and the relative strength of their $\gamma$-ray emission. Here, we compare the properties of relativistic transverse waves observed in two blazars: BL Lac (cohen15; arshakian24) and the FSRQ 3C 279, as analysed in this work. The wave frequency and mean wave amplitude are estimated to be $0.5$ yr-1 and 0.06 mas (0.38 pc) for the 3C 279 and $0.8$ yr-1 and 0.028 mas (0.037 pc) for the BL Lac. While the wave frequencies are comparable, the mean wave amplitude of the 3C 279 is larger than that of the BL Lac by a factor of ten. Assuming that the transverse wave is an Alfvén wave, and the transverse magnetic energy $B_{\perp}$ dominates over the transverse kinetic energy of the jet plasma, the total energy of the wave in the frame of the jet $E_{\rm tot}\propto B_{\perp}^{2}V\approx(A_{\rm w}f_{\rm w}B_{0})^{2}V$, where the magnetic energy of the perturbation $B_{\perp}\approx A_{\rm w}f_{\rm w}B_{0}$, $B_{0}$ is the mean large-scale magnetic field that exists before perturbations by Alfvén waves, and $V$ is the total volume of the jet segment containing the wave. Assuming that the values of $B_{0}$ and $V$ are approximately equal for BL Lac and 3C 279, we obtain the ratio of their total energies $(0.5\times 0.38)^{2}/(0.8\times 0.037)^{2}\approx 40$; that is, the total transverse wave energy of 3C 279 is approximately $10^{1.6}$ times higher than that of the BL Lac wave. This difference is primarily due to the significantly larger mean wave amplitude of the quasar 3C 279.
@@
The mean speeds of the QSCs in 3C 279 and BL Lac are superluminal, $15c$ and $2.5c$ (arshakian24), with the former being about six times higher. This difference can be attributed to a smaller viewing angle of the jet, a higher inclination angle and speed of the transverse wave in 3C 279, or a combination of these factors (see Eq. 1).
@@
Another significant difference lies in the jet structure on pc-scales: the deprojected distance of the QSC from the core in 3C 279 ($\approx 125$ pc) is approximately 40 times greater than that of the QSC in BL Lac ($\approx 3$ pc), assuming a jet viewing angle of $1\degr$ adopted in this work and $6\degr$ from arshakian24, respectively. The QSC in the BL Lac jet is believed to be a recollimation shock at 15 GHz (cohen14), supported by high-resolution observations at 22 GHz with RadioAstron and 43 GHz with VLBA (gomez16). Relativistic magnetohydrodynamics simulations in gomez16 show that recollimation shocks appear at relatively short distances from the core due to the jet’s high speed and a decreasing pressure gradient in the ambient medium, possibly caused by interactions with the broad-line region (0.110.1-1 pc) or torus ($\lesssim 10$ pc). On the other hand, the location of the QSC in 3C 279 at a distance of about 125 pc from the core is beyond the broad-line region and torus, indicating that the recollimation shock is generated by interaction with a hot gas region.
@@
burd22 show that the transition of the jet geometry from parabolic to conical in 3C 279 occurs at a separation of 0.6 mas (deprojected distance of $\approx 210$ pc), likely marking the transition from a magnetically dominated to a particle-dominated jet (kovalev20). In BL Lac, the break point of the jet geometry transition occurs at a separation of 2.5 mas (deprojected distance of $\approx 30$ pc; kovalev20), well beyond the torus region. This supports the dominance of helical magnetic energy over kinetic energy of the jet at large distances from the central black hole, and consequently the presence of relativistic Alfvén waves extending up to the jet geometry transition location.
@@
Lastly, the ratio of rmsc/rmss is $\approx 1.6$ for 3C 279 and $\approx 1.3$ for BL Lac (arshakian20), indicating that the variation in the displacements of the core is more pronounced in 3C 279. This could be attributed to stronger opacity changes (or higher particle density variations) during radio flares or a stronger magnetic field in the core region, as well as geometric effects such as wobbling of the core (see a detailed discussion in arshakian20).